OpenStudy (anonymous):

Would you help me with this integral? ∫ 1 / lnx dx from 3 to + infinity. I need to know it it diverges.

OpenStudy (anonymous):

\[\int\limits_{3}^{\infty} \frac{ 1 }{ \ln x } dx\]

OpenStudy (anonymous):

Wanna work on this together (if you're still around?)

OpenStudy (anonymous):

I just got here.

OpenStudy (anonymous):

I already solved it though. Thanks.

OpenStudy (anonymous):

Awesome. I can tell you it does diverge.

OpenStudy (anonymous):

Ah, great to hear. Have fun!

OpenStudy (anonymous):

Maybe you can help out with my next question.